An open macroeconomic model for a hypothetical economy is represented as follows
Y= C0 +Io+Go+X0-M, M=mo+m1yd,C=co+c1yd, T=tY and Yd=Y-T

a. Show that equal change in tax and government expenditure are expansionary to the economy

b. Government Expenditure = Multiplier

c. Derive the equilibrium level of savings in the economy above
1
Expert's answer
2016-01-19T08:28:41-0500
a. The equal change in tax and government expenditure are expansionary for the economy, because if the government expenditure increases (Go to G1), the GDP will increase too, as Y= C0 +Io+Go+X0-M. So, the equal change in tax and government expenditure will have expansionary effect.
b. Investment multiplier is simply the multiplier effect of an injection of investment into an economy.
The investment multiplier in our case will be mi = 1/(1 - c) = 1/(1 - c1).
c. In the equilibrium savings are equal to investment, so in our case the equilibrium level of savings is S = Io.